Brian M. Carlson Computer Consultant sandals@crustytoothpaste.ath.cx http://crustytoothpaste.ath.cx/consulting 4444 Cullen Boulevard #411 Houston, Texas 77004-2637 United States ☎ +1 832 623 2791 4096R/0223B187:88AC E9B2 9196 305B A994 7552 F1BA 225C 0223 B187 2048R/560553E7:FE82 7C9F EB21 5436 2F96 25BA 927B 0A51 Brian M. Carlson Computer Consultant sandals@crustytoothpaste.ath.cx http://crustytoothpaste.ath.cx/consulting 4444 Cullen Boulevard #411 Houston, Texas 77004-2637 United States ☎ +1 832 623 2791 4096R/0223B187:88AC E9B2 9196 305B A994 7552 F1BA 225C 0223 B187 2048R/560553E7:FE82 7C9F EB21 5436 2F96 25BA 927B 0A51^{Brian M. Carlson Computer Consultant sandals@crustytoothpaste.ath.cx http://crustytoothpaste.ath.cx/consulting 4444 Cullen Boulevard #411 Houston, Texas 77004-2637 United States ☎ +1 832 623 2791 4096R/0223B187:88AC E9B2 9196 305B A994 7552 F1BA 225C 0223 B187 2048R/560553E7:FE82 7C9F EB21 5436 2F96 25BA 927B 0A51 Brian M. Carlson Computer Consultant sandals@crustytoothpaste.ath.cx http://crustytoothpaste.ath.cx/consulting 4444 Cullen Boulevard #411 Houston, Texas 77004-2637 United States ☎ +1 832 623 2791 4096R/0223B187:88AC E9B2 9196 305B A994 7552 F1BA 225C 0223 B187 2048R/560553E7:FE82 7C9F EB21 5436 2F96 25BA 927B 0A51}^{Brian M. Carlson Computer Consultant sandals@crustytoothpaste.ath.cx http://crustytoothpaste.ath.cx/consulting 4444 Cullen Boulevard #411 Houston, Texas 77004-2637 United States ☎ +1 832 623 2791 4096R/0223B187:88AC E9B2 9196 305B A994 7552 F1BA 225C 0223 B187 2048R/560553E7:FE82 7C9F EB21 5436 2F96 25BA 927B 0A51 Brian M. Carlson Computer Consultant sandals@crustytoothpaste.ath.cx http://crustytoothpaste.ath.cx/consulting 4444 Cullen Boulevard #411 Houston, Texas 77004-2637 United States ☎ +1 832 623 2791 4096R/0223B187:88AC E9B2 9196 305B A994 7552 F1BA 225C 0223 B187 2048R/560553E7:FE82 7C9F EB21 5436 2F96 25BA 927B 0A51}^{Brian M. Carlson Computer Consultant sandals@crustytoothpaste.ath.cx http://crustytoothpaste.ath.cx/consulting 4444 Cullen Boulevard #411 Houston, Texas 77004-2637 United States ☎ +1 832 623 2791 4096R/0223B187:88AC E9B2 9196 305B A994 7552 F1BA 225C 0223 B187 2048R/560553E7:FE82 7C9F EB21 5436 2F96 25BA 927B 0A51 Brian M. Carlson Computer Consultant sandals@crustytoothpaste.ath.cx http://crustytoothpaste.ath.cx/consulting 4444 Cullen Boulevard #411 Houston, Texas 77004-2637 United States ☎ +1 832 623 2791 4096R/0223B187:88AC E9B2 9196 305B A994 7552 F1BA 225C 0223 B187 2048R/560553E7:FE82 7C9F EB21 5436 2F96 25BA 927B 0A51}^{Brian M. Carlson Computer Consultant sandals@crustytoothpaste.ath.cx http://crustytoothpaste.ath.cx/consulting 4444 Cullen Boulevard #411 Houston, Texas 77004-2637 United States ☎ +1 832 623 2791 4096R/0223B187:88AC E9B2 9196 305B A994 7552 F1BA 225C 0223 B187 2048R/560553E7:FE82 7C9F EB21 5436 2F96 25BA 927B 0A51 Brian M. Carlson Computer Consultant sandals@crustytoothpaste.ath.cx http://crustytoothpaste.ath.cx/consulting 4444 Cullen Boulevard #411 Houston, Texas 77004-2637 United States ☎ +1 832 623 2791 4096R/0223B187:88AC E9B2 9196 305B A994 7552 F1BA 225C 0223 B187 2048R/560553E7:FE82 7C9F EB21 5436 2F96 25BA 927B 0A51}